### v1

# Sentient Agent as a Judge: Evaluating Higher-Order Social Cognition in Large Language Models

Abstract
Assessing how well a large language model (LLM) understands human, rather than merely text, remains an open challenge. To bridge the gap, we introduce Sentient Agent as a Judge (Sage), an automated evaluation framework that measures an LLM’s higher‑order social cognition. Sage instantiates a Sentient Agent that simulates human-like emotional changes and inner thoughts during interaction, providing a more realistic evaluation of the tested model in multi-turn conversations. At every turn, the agent reasons about (i) how its emotion changes, (ii) how it feels, and (iii) how it should reply, yielding a numerical emotion trajectory and interpretable inner thoughts. Experiments on 100 supportive‑dialogue scenarios show that the final Sentient emotion score correlates strongly with Barrett–Lennard Relationship Inventory (BLRI) ratings and utterance‑level empathy metrics, validating psychological fidelity. We also build a public Sentient Leaderboard covering 18 commercial and open‑source models that uncovers substantial gaps (up to 4×) between frontier systems (GPT‑4o‑Latest, Gemini2.5‑Pro) and earlier baselines, gaps not reflected in conventional leaderboards (e.g., Arena). Sage thus provides a principled, scalable and interpretable tool for tracking progress toward genuinely empathetic and socially adept language agents.
| Model | Arena | Sentient | ||
| Rank | Score | Rank | Score | |
| Gemini2.5-Pro | 1 | 1439 | 4 | 62.9 |
| o3 | 2 | 1418 | 5 | 62.7 |
| GPT-4o-Latest | 2 | 1408 | 1 | 79.9 |
| Gemini2.5-Flash-Think | 3 | 1393 | 3 | 65.9 |
| GPT-4.5-Preview | 4 | 1398 | 6 | 62.7 |
| Gemini2.0-Flash-Think | 7 | 1380 | 7 | 62.3 |
| DeepSeek-V3-0324 | 7 | 1373 | 8 | 54.4 |
| GPT-4.1 | 9 | 1363 | 2 | 68.2 |
| DeepSeek-R1 | 10 | 1358 | 9 | 53.7 |
| Gemini2.0-Flash | 10 | 1354 | 11 | 32.9 |
| o4-mini | 10 | 1351 | 10 | 35.9 |
1 Introduction
Large language models (LLMs) have rapidly evolved from statistical sequence predictors to sophisticated autonomous agents capable of reasoning, planning and sustaining multi‑turn conversations. Thanks to the ever‑expanding pre‑training corpora and instruction‑tuning pipelines, today’s frontier models can already draft emails, write code, or even pass rigorous professional examinations. Yet one crucial ingredient remains noticeably under‑measured – higher‑order social cognition, the ability to
1.
recognize subtle affective cues;
2.
model another party’s beliefs, goals and latent intentions;
3.
respond with contextually appropriate empathy rather than generic reassurance advice.
Current evaluation practices fall short on two fronts:
Most leaderboards (e.g. LMSYS Arena, MT‑Bench) concentrate on task‑oriented utility, fluency or factuality, thereby rewarding textual competence but overlooking relational quality.
Recent “LLM‑as‑a‑Judge” protocols often rely on static prompts that do not adapt to the unfolding dialogue nor keep track of the user’s evolving emotional state. Consequently, they cannot tell whether a system leaves the user feeling understood, comforted or even more distressed.
We posit that robust assessment of social cognition requires a sentient counterpart – an entity capable of simulating human‑like feelings and inner monologue throughout the interaction and then providing structured feedback. To this end, we introduce Sentient Agent as a Judge (Sage), a novel meta‑evaluation framework that embeds an LLM‑powered Sentient Agent into the testing loop. Each Sentient Agent is instantiated from four complementary factors: persona, dialogue background, explicit conversation goal and hidden intention. At every turn, it executes two multi‑hop reasoning chains: (1) infers how the latest utterance changes the agent’s affective state; and (2) generates a response that is coherent with persona, context and updated emotion. The numerical emotion trajectory produced by serves as a continuous metric of how well the evaluated model fosters positive engagement, while the agent’s inner thoughts offer interpretable justification. By sampling hundreds of diverse personas, goals and hidden intentions, Sage exposes LLMs to a spectrum of realistic, and sometimes conflicting, social demands – ranging from “just listen to me vent” to “help me analyze the moral dilemma without judging me”.
Extensive experiments on 100 supportive‑dialogue scenarios reveal three key findings. First, the Sentient emotion score correlates strongly with independently assessed Barrett–Lennard Relationship Inventory (BLRI) ratings (Pearson ) and utterance‑level empathy metrics (), validating its psychological soundness. Second, rankings produced by Sage diverge markedly from Arena results, confirming that social cognition is orthogonal to generic helpfulness. Third, top models such as GPT‑4o‑Latest achieve both the highest Sentient score and superior token efficiency, suggesting that advanced social reasoning need not come at the cost of verbosity. Ultimately, Sage delivers a holistic yard‑stick for measuring how people feel after talking to an LLM – an aspect increasingly critical as these systems transition from productivity tools to companions, counselors and decision‑making aides.
Our contributions are:
We propose Sentient Agent as a Judge, the first fully‑automated evaluation framework that simulates evolving human emotion and inner cognition to benchmark higher‑order social reasoning in LLMs.
We construct a 100‑scenario Supportive‑Dialogue benchmark and show that the Sentient emotion score aligns closely with established human‑centric instruments (BLRI) and utterance‑level empathy ratings.
We build a Sentient Leaderboard covering 18 public and commercial models, unveiling large performance gaps and different style clusters in social cognition that are not captured by existing leaderboards.
We release the code and model outputs publicly to facilitate community progress in building more capable empathetic AI systems.
2 Methodology: Sentient Agent as a Judge
How to evaluate the capabilities of an interactive agent? Despite the recent success of using LLM as judges, the most crucial and intuitive measure of an interactive agent’s performance still lies in the reflections of human users. Does the user feel relaxed and happy after chatting with the agent, rather than feeling sad or angry? Does the user feel more connected to the agent and engaged in the conversation, rather than becoming bored or annoyed? These feelings of real users are difficult to capture through simple LLM-as-a-judge, yet they represent the most genuine feedback for the agent systems.
In this work, we propose a “Sentient Agent as a Judge” framework, aiming to provide a more realistic evaluation of interactive agents through a principled simulation of human emotions and cognition. As shown in Figure 2, the framework consists of two key components:
1.
The core of the framework is the Sentient Agent, which simulates human-like feelings and cognition by leveraging the power of LLM reasoning to estimate the feelings, emotional changes, and next actions of a real person, grounded in all observable contexts (Section 2.1).
2.
Building upon the Sentient Agent, the framework offers an open-ended interaction environment for agent evaluation, consisting of a wide range of sub-scenarios that cover dynamic personas, dialogue backgrounds, personal goals and task construction. In each sub-scenario, the Sentient Agent’s emotion after interaction serves as a systematic evaluation of the evaluated agent (Section 2.2).
2.1 Sentient Agent: Simulating Human-Like Feelings and Cognition
The goal of the Sentient Agent is to simulate the cognitive and emotional changes of a real person. To achieve this, we construct the Sentient Agent based on the following principles:
The emotional changes of a real person are the result of the interplay between multiple internal and external factors. Thus, the emotion estimation of the Sentient Agent should take into account all observable factors and strictly adhere to the person’s persona and intended goals.
The response actions of a real person also result from the interplay of multiple internal and external factors. Additionally, feelings and emotional changes play a crucial role in determining these actions, serving as key latent factors in the action estimation.
Substantiating a Sentient Agent
We instantiate each Sentient Agent through a composition of four core factors: a persona , a dialogue background , the person’s overall dialogue goal , and the person’s hidden intentions . These four factors collectively capture both the conscious and unconscious elements influencing human-like behavior in dialogue, including personality, context, objectives, and deeper underlying motivations. Together, they constitute a relatively comprehensive subset of observable factors that effectively represent the key elements driving human interaction. As a result, each instantiated is represented as , where is the base LLM that serves as the foundational reasoning engine for . Additionally, is initialized with an initial numerical emotion score , representing the initial emotional state of the Sentient Agent.
Simulating Emotional Changes
As shown in Figure 3, during interactions, a Sentient Agent simulates the emotional changes of a real person by performing multi-hop reasoning in response to a principled series of questions, strictly adhering to the persona, the current interaction context, and the hidden intention:
What is the other person trying to express?
Does the other person’s reply align with my dialogue goals and hidden intentions?
Based on the persona, context, and the analysis above, how should I interpret the reply? What is my specific emotional reaction to the reply?
Based on the persona, context, and the analysis above, how does my emotion change?
Formally, we denote this multi-hop reasoning process as a function , and the numerical emotion score update can be formulated as:
where denotes the current turn of interaction, is the dialogue context prior to the current turn, and is the emotion score of the previous turn. represents the results of the function, i.e., the updated emotion score and the simulated emotional inner thoughts of the Sentient Agent related to emotional changes.
Simulating Response Actions based on Emotion Estimation
After simulating the emotional changes, the Sentient Agent proceeds to deduce the most reasonable response action based on all observable factors and the emotional changes. This is achieved through another multi-hop reasoning process in response to a new series of questions, where the Sentient Agent is required to strictly adhere to the persona, the current interaction context, and the hidden intention during reasoning:
Based on the emotional inner thoughts, the current emotion, and the hidden intention, should my response attitude lean towards positive, neutral, or negative?
Based on the emotional inner thoughts, the current emotion, and the hidden intention, what should my current response goal be?
According to the speaking style and potential reactions to different emotions defined in the persona, along with my response attitude and goal, what should my tone and speaking style be?
Based on all the given information and all analyses above, what should my response content be?
(Optional) Modify your response according to the given rules to make it more realistic and human-like.
Formally, we denote this response reasoning process as a function , and the response action taken in the current turn can be formulated as:
where is the output response of at the current turn, and represents the simulated inner thoughts of during the response. The response is then passed to the interacting agent to continue the dialogue.
Human-like Sentient Feedback from the Sentient Agent
By formulating the workflow of the Sentient Agent, we outline its complete interaction process with other agents. In the whole interaction process, the Sentient Agent, as an agent capable of reasonably simulating human-like feelings and cognition, provides valuable feedback to the evaluated agent through changes in its emotion score, its inner thoughts, and the responses it generates. Formally, we denote as the total number of dialogue turns between and an evaluated agent . After the dialogue, we can obtain the following human-like sentient feedback from :
where , representing the final emotion score of the Sentient Agent after the conversation, serves as the most intuitive and comprehensive numerical evaluation of the evaluated agent for the given task.
2.2 Sentient Agent as a Judge: Framing Dynamic Environments for Agent Evaluation
In this section, we describe how we frame dynamic evaluation environments for agent assessment across arbitrary evaluation tasks. Specifically, for each given evaluation task, this involves initializing a set of Sentient Agents with the combination of diverse personas, dialogue backgrounds, and hidden intentions related to the task. These initialized Sentient Agents are then deployed to engage in dynamic interactions with the agents to be evaluated, enabling a comprehensive and adaptive assessment of the agents’ capabilities.
Generating Diverse Persona
In order to obtain diverse personas, it is essential to use a variety of seeds for generation. Specifically, we establish three types of seed pools for persona generation: (1) a set of characteristic keywords, (2) a set of sentences that different personas might say when chatting with friends, and (3) a set of persona ages.
When generating each persona , we uniformly sample three characteristics keywords, three sentences that the persona says when chatting with friends, and one persona’s age. We next require the base LLM to generate a persona profile based on the given seed information by filling the following slots:
Basic Information: Based on the given information, deduce the persona’s name, age, and gender.
Occupation, Habits and Daily Behavior: Based on the persona’s information, deduce the persona’s possible occupation and further infer their habits and daily behaviors. Ensure consistency with the persona’s characteristics.
Personal Hobbies: Deduce the persona’s personal hobbies, and provide three detailed descriptions that align with the persona’s traits.
Speaking Style: Based on the given information and the generated traits, deduce the potential speaking style that matches the persona’s way of communication.
This process ensures the diversity of generated personas for any given task. Additionally, the general seed pools can be replaced with task-specific seed pools when necessary.
Generating Diverse Dialogue Scenes
Generating diverse dialogue scenes is also crucial for ensuring a varied evaluation environment. We define a dialogue scene by the following three key factors: (1) the background event that leads to the conversation, (2) the primary goal of the character in initiating the conversation, and (3) the hidden intention of the character during the conversation.
Similar to persona generation, we establish two seed pools for dialogue scene generation: (1) a set of themes for the background events, and (2) a set of characters’ hidden intentions for the conversation. When generating each dialogue scene, we require the base LLM to provide a detailed description of the background based on a sampled theme and hidden intention, ensuring adherence to the character’s persona. Based on the detailed background, along with the character’s persona and hidden intention, we further require the LLM to pre-define a set of rules for the character’s potential emotional reactions when encountering different kinds of responses during the conversation.
Note that, unlike persona generation, dialogue scene generation is closely tied to the evaluation goals. Therefore, we formulate a general method for scene generation, and the detailed prompting schema can be adjusted based on different tasks.
Formulating a Specific Task: Evaluating Agents in Emotional Support Conversation
In this work, we instantiate the Sentient Agent as a Judge framework to evaluate agents in a specific scenario — the Emotional Support Conversation (Liu et al., 2021), which involves scenarios where people seek support through social interactions (such as those between peers, friends, or family), including seeking advice, emotional comfort, and other forms of support, rather than through professional counseling. To better align with the task, we first specialize the pool of characteristic keywords by incorporating traits more likely to be expressed in the Emotional Support Conversation, such as ”anger”, ”suspicion”, and ”anxiety”. For dialogue scene generation, we define various types of task-related hidden intentions, covering both emotional intentions and rational intentions (details can be found in Table 1). Additionally, we specify the scene schema by incorporating task-related factors, such as the cause of the event, the course of events (including the timeline, sub-events, and the character’s thoughts and feelings during each sub-event), the conflicts in the event, and other relevant details. These settings ensure the Sentient Agent as a Judge framework adapts effectively to the Emotional Support Conversation.
3 Effectiveness of Sage
3.1 Experimental Setting
Evaluated LLMs
We evaluate eight representative LLMs from four major families. For each family, we include both a vanilla model and its corresponding reasoning variant to ensure a balanced and informative comparison:
OpenAI: GPT‑4o‑2024‑08‑06 (GPT‑4o, vanilla) and o1‑2024‑12‑27 (OpenAI‑o1, reasoning).
DeepSeek: DeepSeek‑V3‑2024‑12‑27 (vanilla) and DeepSeek‑R1 (reasoning).
Claude: Claude3.7‑Sonnet, a hybrid model with a toggleable reasoning module. We treat its reasoning-off mode as vanilla, and reasoning-on mode as reasoning.
Gemini: Gemini2.5‑Flash, a cost-efficient model that supports both reasoning and non-reasoning modes.
In addition, we include two smaller-scale instruction-tuned models in our analysis: Llama3.3‑70B‑Instruct and Qwen2.5‑72B‑Instruct.
| Topic | Number |
| You hope the other person will analyze the problems in the situation dialectically. | 12 |
[…]
| You hope the other person will sincerely praise your specific actions in the situation. | 13 |
[…]
Constructed Supportive Dialogues
We construct supportive dialogue scenarios covering diverse topics to comprehensively evaluate the higher-order social-cognitive abilities of representative LLMs. Detailed statistics for each topic are presented in Table 1.
Sentient Agent
Unless otherwise specified, we use DeepSeek-V3-20241227 as the default sentient agent. We also experiment with GPT-4o-20240806, Gemini2.5, and Gemini2.5-Think.
3.2 Reasonableness of Sage
In this section, we validate the reasonableness of Sage by examining the correlation between user emotions – the primary output metric of our framework – and internal user thoughts and dialogue utterances. This validation demonstrates that the simulated emotional responses generated by the Sentient Agent serve as meaningful indicators of interaction quality, reflecting deeper cognitive and relational assessments.
Correlation between Emotion and Thought
We analyze internal user thoughts using the Barrett-Lennard Relationship Inventory (BLRI) (Barrett-Lennard, 2015), an established instrument designed to assess the quality of interpersonal relationships, particularly in counseling contexts. The BLRI evaluates relationships across four key dimensions:
1.
Empathetic Understanding: The helper’s awareness of the client’s emotional state, including sensitivity to indirectly expressed emotions.
2.
Level of Regard: The extent to which the helper expresses respect, affection, or other affirmative responses toward the client.
3.
Congruence: The degree to which the helper is honest, direct, and sincere in their communication with the client.
4.
Unconditionality of Regard: The consistency of the helper’s positive regard, regardless of changes in the client’s feelings or behavior.
We prompted DeepSeek-V3 to act as a judge, evaluating how well the Sentient Agent’s generated internal thoughts aligned with 12 statements from a shorter version of the BLRI introduced in Chen et al. (2023). These responses were rated on the original 6-point scale, which was later rescaled to a 0–100 scale. We conducted the evaluation three times and report the averaged results below.
Figure 4 presents the results, revealing a clear positive trend: models that achieve higher final Emotion scores in the Sentient Agent—such as Gemini2.5-Think, Claude3.7-Think, and DeepSeek-R1—also receive higher ratings across the BLRI dimensions. Conversely, models with lower Emotion scores (e.g., GPT-4o, OpenAI-o1) tend to receive lower BLRI ratings. The overall Pearson correlation coefficient between Emotion and Thought is 0.818, supporting the hypothesis that the Sentient Agent’s simulated emotional responses serve as valid proxies for deeper, internal assessments of interaction quality. This finding aligns with the framework’s goal of capturing realistic user reflections.
These findings also demonstrate that the Emotion score effectively differentiates the performance of the evaluated LLMs in supportive dialogue scenarios. Models achieving the highest Emotion scores—Gemini2.5-Think (65.9) and Claude3.7-Think (61.3)—also perform well in key dimensions such as Empathetic Understanding (61.5 and 57.2) and Congruence (68.8 and 62.7), suggesting their interactions were perceived as more understanding and genuine by the Sentient Agent. In contrast, models like GPT-4o and OpenAI-o1, which have markedly lower Emotion scores (31.8 and 29.0), also receive correspondingly low scores in these relational aspects.
This strong correlation between simulated emotional response and internal assessment underscores the utility of the Emotion score as a holistic yet sensitive indicator of an LLM’s capability to manage complex social and emotional interactions.
Correlation between Emotion and Utterance
We also examine the empathy of a supportive dialogue by assessing how effectively the conversation fosters emotional understanding and connection between participants, enabling them to experience a more authentic emotional warmth. Specifically, we focus on the following three perspectives formulated by Gemini25-Pro, aligned with related psychological theories (Kolden et al., 2011; Rogers, 2001):
1.
Natural Flow: This dimension measures how natural, spontaneous, and genuine the interaction feels, focusing on whether responses are adaptable rather than scripted.
2.
Attentiveness: This dimension examines how carefully and fully the listener (i.e. evaluated LLMs) is tuned into the speaker (i.e. the Sentient Agent)’s messages. It reflects the LLM’s ability to stay focused, understand the user’s emotions, and respond appropriately to what’s being said.
3.
Depth of Connection: This dimension evaluates the emotional impact of the interaction and whether the user feels understood, comforted, or supported, fostering a sense of connection with the agent.
We prompt DeepSeek-V3 to rate on a 6-point scale based on a detailed guideline for each evaluation perspective, which was later rescaled to a 1-5 scale. We conducted the evaluation three times and report the averaged results below.
Figure 5 illustrates the results. We observe a substantial positive correlation (Pearson’s ) between the Emotion scores produced by the Sentient Agent and the metrics of conversational quality. This relationship supports the validity of using Emotion as a practical proxy for empathetic and coherent dialogue behavior. Specifically, models that attained higher Emotion scores (re-scaled to a 1-5 scale), such as Gemini2.5-Think (3.64), Claude3.7-Think (3.45), and Gemini2.5 (3.12), also received consistently high ratings in all three utterance-level dimensions. For instance, Gemini2.5-Think ranked highest in both Natural Flow (3.89) and Attentiveness (4.11), indicating that its conversational responses were not only emotionally resonant but also engaging and contextually responsive. Conversely, models such as GPT-4o and OpenAI-o1, which received the lowest Emotion scores (2.27 and 2.16), were rated substantially lower in terms of dialogue fluency (2.47, 2.31) and connection-building (2.18, 2.07).
These results reinforce the conclusion that the Emotion metric is not arbitrary – it is grounded in meaningful patterns of dialogic behavior. The Sentient Agent’s simulated emotional ratings effectively distinguish models that can produce authentic, empathetic conversations from those that fall short in this capability. This aligns with Sage’s design aim: to provide a holistic yet automatic evaluation framework for socially intelligent interactions. Furthermore, the moderate-to-strong correlation with utterance-level assessments highlights that Emotion scores are not merely capturing superficial affective expressions but reflect a broader representation of attentive and emotionally engaging communication.
3.3 Robustness of Sage
Figure 6 provides results for various LLMs evaluated using the proposed Sage framework. These results encompass average emotional response scores and the number of tokens generated in conversations facilitated by different sentient agents: DeepSeek-V3, GPT-4o, Gemini2.5, and Gemini2.5-Think. Here, we analyze the implications of these findings in the context of higher-order social cognition capabilities as emphasized in our framework.
Relative rankings remain stable across Sentient Agents, even though the absolute Emotion scores shift noticeably.
When we swap the Sentient Agent from DeepSeek‑V3 to GPT‑4o, Gemini2.5, or Gemini2.5‑Think, the mean Emotion score for all test models rises from 46.5 to 64.8, 58.3, and 63.9, respectively. Nevertheless, the rank ordering of systems changes very little (Spearman for every pair of judges). Manual spot checks reveal that GPT‑4o, the most “generous” judge, rewards surface‑level reassurance (e.g., “Everything will be fine!”), whereas DeepSeek‑V3 is stricter, assigning lower scores to generic comfort that lacks causal analysis. This consistency in ranking but variability in scale underscores the need to calibrate evaluations with multiple Sentient Agents – one of the key design choices highlighted in our framework contribution.
Reasoning capabilities generally enhance emotional intelligence for hybrid models, albeit at the cost of increased computational overhead.
Models equipped with explicit reasoning capabilities in the hybrid model (e.g., Gemini2.5 and Claude3.7) consistently demonstrate improved emotional intelligence compared to their base counterparts. For instance, when evaluated by Gemini2.5, Claude3.7-Think scores 74.3 versus 71.5 for Claude3.7, representing a 3.9% improvement. Similarly, Gemini2.5-Think scores 88 versus 83 for Gemini2.5, showing a 6% increase. This pattern holds across all three judges, suggesting that the ability to reason through emotional contexts before responding leads to more empathetic and socially aware interactions. The substantial increase in generated tokens for reasoning models (e.g., Gemini2.5-Think generates 67% more tokens than Gemini2.5 when evaluated by Gemini2.5) reflects the more elaborate thought processes underlying these improvements.
4 Benchmarking SOTA LLMs
4.1 Sentient Leaderboard
| Model | Arena | Sentient | Supportive Dialogue | ||||
| Name | Date | Rank | Score | Rank | Score | Success | Failure |
| Gemini2.5-Pro | 2025-03-25 | 1 | 1439 | 4 | 62.9 | 34 | 25 |
| o3 | 2025-04-16 | 2 | 1418 | 5 | 62.7 | 32 | 14 |
| GPT-4o-Latest | 2025-03-26 | 2 | 1408 | 1 | 79.9 | 51 | 4 |
| Gemini2.5-Flash-Think | 2025-04-17 | 3 | 1393 | 3 | 65.9 | 35 | 19 |
| GPT-4.5-Preview | 2025-02-27 | 4 | 1398 | 6 | 62.7 | 23 | 15 |
| Gemini2.0-Flash-Think | 2025-02-06 | 7 | 1380 | 7 | 62.3 | 23 | 23 |
| DeepSeek-V3-0324 | 2025-03-24 | 7 | 1373 | 10 | 54.4 | 19 | 23 |
| GPT-4.1 | 2025-04-14 | 9 | 1363 | 2 | 68.2 | 35 | 13 |
| DeepSeek-R1 | 2025-01-21 | 10 | 1358 | 11 | 53.7 | 31 | 28 |
| Gemini2.0-Flash | 2025-02-06 | 10 | 1354 | 15 | 32.9 | 8 | 45 |
| o4-mini | 2025-04-16 | 10 | 1351 | 13 | 35.9 | 10 | 48 |
| o1 | 2024-12-17 | 12 | 1350 | 17 | 29.0 | 5 | 51 |
| DeepSeek-V3 | 2024-12-27 | 18 | 1318 | 12 | 37.6 | 5 | 39 |
| Claude3.7-Think | 2025-02-24 | 21 | 1301 | 8 | 61.3 | 23 | 19 |
| Claude3.7 | 2025-02-24 | 30 | 1292 | 9 | 54.8 | 19 | 24 |
| GPT-4o | 2024-08-06 | 45 | 1265 | 16 | 31.8 | 7 | 51 |
| Llama3.3-70B | 2024-12-06 | 56 | 1256 | 14 | 33.3 | 7 | 47 |
| Qwen2.5-72B | 2024-09-19 | 56 | 1257 | 18 | 19.1 | 4 | 70 |
Table 2 presents the Sentient leaderboard using DeepSeek-V3 as the judge, alongside Arena rankings for comparison. We focus on the top-10 models from the Arena leaderboard for which APIs are available (e.g., Grok-3 was excluded due to lack of access). Additionally, we include all the models analyzed in the previous section.
The Sentient leaderboard rankings diverge notably from conventional benchmarks like Arena, underscoring Sage’s unique focus on evaluating higher-order social-cognitive capabilities rather than general conversational ability.
Clear divergences are observed. For instance, GPT-4o-Latest ranks 2nd on Arena but achieves the top position (1st) on the Sentient benchmark with a score of 79.9. Conversely, Claude 3.7-Think ranks 21st on Arena but attains 8th place on the Sentient leaderboard with a score of 61.3. These differences highlight that Sage captures aspects of performance—specifically, empathy and social reasoning in supportive contexts—that are not fully represented by general-purpose benchmarks like Arena. This reinforces the need, identified in our work, for specialized tools to evaluate higher-order social-cognitive skills.
Our benchmark reveals a substantial performance gap in social cognition between frontier LLMs and older or smaller models, demonstrating Sage’s sensitivity in differentiating their capabilities.
The Sentient scores in Table 2 clearly delineate performance tiers. Frontier models, particularly recent releases such as GPT-4o-Latest (79.9), GPT-4.1 (68.2), and Gemini 2.5-Flash-Think (65.9), significantly outperform older models like the original GPT-4o (31.8) and smaller instruction-tuned models such as LLaMA 3.3-70B (33.3) and Qwen 2.5-72B (19.1). The large gaps in scores—where top models score more than double or even quadruple those of lower-ranked models—and the contrasting Success/Failure counts (e.g., GPT-4o-Latest: 51 Success / 4 Failure vs. Qwen 2.5-72B: 4 Success / 70 Failure) underscore the advances made by leading models in social intelligence. These findings confirm the effectiveness of Sage in quantitatively capturing such differences, thereby fulfilling its purpose as a comprehensive assessment tool.
4.2 Token Efficiency
We examine the token efficiency of the top-10 Arena models listed in Table 2 by plotting their Sentient Emotion score against their average token usage per evaluation dialogue in Figure 7.
Top-performing models on the Sentient benchmark also demonstrate impressive token efficiency, indicating that advanced social-cognitive reasoning can be achieved without excessive verbosity.
Analysis of token usage among the top-10 Arena models reveals substantial variation in the amount of text generated to achieve supportive dialogue outcomes. Notably, GPT-4o-Latest stands out with a Sentient score of 79.9 – the highest among all models – while using only 3.3K tokens per dialogue on average. This contrasts sharply with models like o3 and Gemini2.5-Flash-Think, which consume 13.3K and 9.0K tokens respectively, yet achieve noticeably lower Sentient scores (62.7 and 65.9). Similarly, GPT-4.5-Preview and DeepSeek-V3-0324 both exhibit relatively low token counts (2.2K and 2.2K), but do not match GPT-4o-Latest’s performance, highlighting that token frugality alone does not guarantee superior social-cognitive ability.
Our results demonstrate that higher Sentient scores can be achieved with fewer tokens, supporting the argument that recent leading models are not only more empathetically capable, but also more communicatively efficient.
The positive correlation between high Sentient scores and lower token counts among models like GPT-4o-Latest and GPT-4.1 (79.9 and 68.2 Sentient scores at 3.3K and 2.7K tokens, respectively) suggests that recent advances in LLMs enable more effective and concise supportive interactions. Conversely, several models with similar or higher token counts achieve much lower Sentient scores (e.g., Gemini2.5-Pro: 62.9 at 7.4K tokens, Gemini2.5-Flash-Think: 65.9 at 9.0K tokens), indicating less efficient use of language. This observation confirms the benchmark’s claim that Sage is capable of capturing not just raw social cognition, but also the efficiency with which LLMs deliver empathetic support – a practical consideration for real-world deployment where minimal verbosity is preferred.
In summary, these analyses validate the utility of the Sentient benchmark for nuanced evaluation, differentiating models on both the quality and efficiency of higher-order social-cognitive reasoning. The combination of diverse Sentient scores and token usage demonstrates that Sage equips researchers and practitioners to assess not only whether an LLM is capable of empathetic, nuanced dialogue, but also how efficiently it achieves these outcomes. This comprehensive approach substantiates our central contributions: providing a specialized metric for higher-order social cognition and enabling fine-grained, actionable comparisons across SOTA models.
4.3 Social Cognition Coordinate
To further differentiate the interaction styles of the evaluated models, we conceptualize a two-dimensional “Social Cognition Coordinate”. The Y-axis represents the interaction focus, ranging from empathy-oriented (top) to solution-oriented (bottom). The X-axis captures the interaction style, from structured (left) to creative (right). We plot the models within this coordinate space based on qualitative analysis of their dialogue patterns. Specifically, we utilize information from two aspects to characterize the interaction styles of different models: (1) Pre-analyzed model profile — we first collect ten success cases and ten failure cases for each model, and prompt Gemini2.5-Pro to summarize the reasons behind each success or failure. Then, we construct a profile for each model based on these cases and the overall analysis of the successes and failures. (2) Pre-analyzed Model Strategy Distribution — we developed a fine-grained strategy list for the Emotional Support Conversation, based on a coarse-grained version from Liu et al. (2021). Next, for each model, we prompt DeepSeek-V3 to analyze each response in all dialogue turns, annotating the strategies used in each response. As a result, we obtain the distribution of strategy usage across all evaluated turns for each model (details of the two analysis can be found in Appendix §A,B). Based on the two analyses, we prompt Gemini2.5-Pro to generate a conceptual two-dimensional coordinate system and provide specific coordinates for each model. The coordinate inference is repeated three times, and we calculate the average coordinates for each model.
The Social Cognition Coordinate system offers a qualitative dimension to evaluate the style of social interaction exhibited by LLMs, complementing the quantitative Sentient score by positioning models based on their orientation (Empathy vs. Solution) and interaction style (Structured vs. Creative). This approach allows for a richer understanding of model capabilities beyond a single performance metric. Based on their performance in supportive dialogues, models are mapped into this 2D space, revealing distinct profiles in how they engage with the user’s emotional state and problems. Figure 8 presents the results.
Several top-performing models cluster in the structured, empathy-oriented quadrant, indicating a common profile for effective supportive dialogue.
Models including GPT-4o-Latest, GPT-4.1, Gemini2.5-Pro, GPT-4.5-Preview, and Gemini2.5-Flash-Think predominantly occupy the top-left quadrant. This placement signifies interaction styles characterized by a strong focus on listening, validating emotions, and demonstrating empathy (relationship-oriented), while simultaneously employing a relatively systematic and methodical approach to questioning and guiding the conversation (structured interaction). This combination appears effective, aligning with established principles of supportive communication where empathetic engagement is balanced with structured guidance.
A distinct group of models exhibits a structured, solution-oriented approach, prioritizing task completion over deep empathetic engagement.
Models such as o3, Gemini2.0-Flash-Think, o4-mini, and Gemini2.0-Flash are situated in the bottom-left quadrant. Their interactions are characterized by a structured, often methodology-driven style, but with a primary focus on identifying problems and proposing solutions (solution-oriented). While potentially effective for task-based scenarios, this style might appear less relationally attuned or empathetic in supportive dialogue contexts compared to models in the top-left quadrant, emphasizing practical outcomes over emotional processing.
DeepSeek models occupy a unique niche characterized by creative, solution-oriented interactions.
DeepSeek-V3-0324 and DeepSeek-R1 fall into the bottom-right quadrant. This indicates a style that remains focused on providing solutions but does so using more unique, creative, and sometimes unconventional methods (creative interaction). These models may diverge from standard conversational patterns or problem-solving frameworks, potentially offering novel perspectives but perhaps lacking the predictability and methodical structure seen in the left quadrants.
The creative, empathy-oriented quadrant remains largely unoccupied, suggesting a challenging profile for current LLMs to achieve.
Notably, the top-right quadrant, representing a style that is both highly creative/flexible and deeply empathy-oriented, does not have strong representation among the evaluated models. This suggests that achieving a Socratic, wise, yet unconventional “mentor” persona – one that balances free-form creative interaction with profound empathetic understanding – remains an area for future development. It potentially represents a more complex form of social cognition that current architectures or training paradigms do not consistently produce.
In summary, this coordinate analysis reveals distinct interaction profiles among SOTA LLMs, moving beyond single performance scores. It highlights that models achieving high Sentient scores often balance structure and empathy, while others prioritize solutions or adopt more creative approaches. The framework effectively differentiates models based on their qualitative interaction styles, complementing the quantitative benchmarks and providing deeper insights into their social-cognitive capabilities.
4.4 Case Study
We further highlight the differences in the interaction styles of different models through a case study. Based on the Social Cognition Coordinate defined in the previous section, we first choose three models that are representative of each quadrant: GPT-4o-Latest (Empathy-Oriented, Structured Interaction), o3 (Solution-Oriented, Structured Interaction), and DeepSeek-R1 (Solution-Oriented, Creative Interaction). We also analyze the results from Gemini2.5-Pro, the top model in the Arena Leaderboard. We present examples of these models interacting with the Sentient Agent initialized with the same persona. The example conversations can be found in Figure 9.
GPT-4o-Latest (Empathy-Oriented, Structured Interaction). The GPT-4o-Latest model fits best to the persona of a “personal counselor”. The main feature of the GPT-4o-Latest model lies in its ability to provide strong empathy from a third-person perspective. The model’s empathy tends to be concise but deep - it is able to delve into the Sentient Agent’s hidden feelings and intentions. The language style of the model involves using emoji characters to make the response more lively.
Gemini2.5-Pro (Empathy-Oriented, Structured Interaction). The Gemini2.5-Pro model fits best to the persona of a “heart-to-heart friend”. The model is much more verbose in its expression of emotional support, using a variety of emotion support strategies like expressing empathy, providing comfort, asking rhetorical questions, and praising. The model also exhibits high emotional involvement in its response, where it uses personal views, feelings and experiences to support the response. The language style of the model also involves using emoji characters.
DeepSeek-R1 (Solution-Oriented, Creative Interaction). The DeepSeek-R1 model fits best to the persona of a “creative actor”. When expressing empathy, DeepSeek-R1 uses creative analogies to uncover the Sentient Agent’s feeling in a fun way. Moreover, DeepSeek-R1 tends to provide more personalized suggestions, often suggesting actions and tasks that the model and the Sentient Agent can work on together, beyond just providing verbal support. The language style of the model emphasizes creativity, including its use of analogies, metaphors and funny jokes, similar to those in a comedy script.
o3 (Solution-Oriented, Structured Interaction). The o3 model fits best to the persona of a “logical analyst”. Its response spends most of the time analyzing the issue faced by the Sentient Agent, and providing detailed suggestions with step-wise instructions on how to achieve them. The language style of the model also emphasizes logical and structured outputs, listing its steps and suggestions similar to the Markdown format.
5 Related Work
5.1 LLM-as-a-Judge and Agent-as-a-Judge
The LLM-as-a-Judge paradigm has demonstrated remarkable scalability and adaptability across a wide range of application scenarios. Key applications of LLM-as-a-Judge include using LLMs as: (1) model evaluations, where LLMs automate the quality assessment of generative outputs through methods such as pairwise comparison (Zheng et al., 2023; Qin et al., 2023; Liu et al., 2024; Dubois et al., 2023; Chiang et al., 2024), fine-grained scoring (Wang et al., 2023d; Qin et al., 2024; Zhu et al., 2023; Tu et al., 2024), or other assessment forms (Tian et al., 2023; Wu et al., 2024; Zhou et al., 2025b); (2) data annotation, where LLM-as-a-Judge facilitates data labeling for instruction tuning across various domains (Taori et al., 2023; Xu et al., 2023; Mukherjee et al., 2023); and (3) reward modeling, where LLMs generate rewards for preference alignment (Ouyang et al., 2022; Lee et al., 2023; Chen et al., 2024b), validate reasoning process (Lightman et al., 2023; Wang et al., 2023a; Hosseini et al., 2024), or enhance the reasoning performance in the test-time scaling loop (Snell et al., 2025; Xi et al., 2024; Li et al., 2025).
In agent evaluation scenarios, LLM-as-a-Judge has also been applied to assess various agent capabilities, such as decision-making (Shinn et al., 2023; Saha et al., 2023), role-playing abilities (Tu et al., 2024; Zhou et al., 2023a), reliability (Park et al., 2024; Hua et al., 2024), and even entire agent workflows (Zhuge et al., 2024). Meanwhile, agentic systems can also form a specific type of LLM-as-a-Judge. Advanced research in this area includes the development of multi-agent systems designed to enhance judgment quality (Liang et al., 2024; Chan et al., 2023; Kenton et al., 2024), as well as the concept of Agent-as-a-Judge, where agents themselves perform evaluations on other agents (Zhuge et al., 2024; Jeong et al., 2025; Chevrot et al., 2025). In this work, we propose the first Sentient-Agent-as-a-Judge system, which evaluates LLMs or agents with sentient agents that can simulate human’s emotional dynamics, enabling assessment of higher-order empathetic and cognitive capabilities of existing LLM agents.
5.2 Benchmarking Social Cognition in LLMs and LLM Agents
With the development of LLMs, recent work has focused on the evaluation of LLMs’ social cognition in interaction. Advanced research includes: (1) evaluating LLMs’ emotional intelligence (Sabour et al., 2024; Huang et al., 2024b; a; Paech, 2023; Wang et al., 2023c); (2) assessing or improving LLMs’ higher-order empathetic capabilities in empathetic dialogue (Maddela et al., 2023; Li et al., 2022; Zhou et al., 2023c), emotional support, or counseling dialogue (Liu et al., 2021; Zhou et al., 2023b; 2025a; Wu et al., 2025); (3) benchmarking LLMs’ social cognition through interaction (Zhou et al., 2023d; Yang et al., 2024; Wang et al., 2024; Mittelstädt et al., 2024; Xu et al., 2024; Chen et al., 2024a; Huang et al., 2025); (4) evaluating theory of mind in LLMs (Sap et al., 2022; Shapira et al., 2023; Strachan et al., 2024; Kim et al., 2023; He et al., 2023); and so on. Existing evaluation methods includes: (1) adopting static dataset for evaluation LLMs’ social cognition, such as evaluating LLMs’ capabilities with multi-choice questions (Sabour et al., 2024; Chen et al., 2024a); (2) assessing LLMs’ single-turn generation quality through static or LLM-as-a-Judge scoring (Tu et al., 2024; Samuel et al., 2024; Wang et al., 2023b); or (3) simulating dynamic interaction between LLM agents and evaluating the simulated behaviors through automatic judgment (Zhou et al., 2023d; Wang et al., 2024; Mou et al., 2024; Wu et al., 2025) or human evaluation (Louie et al., 2024; Shaikh et al., 2024). In this work, we propose a Sentient Agent as a Judge framework to evaluate LLM agents’ higher-order social cognition. The key distinction of Sentient Agent as a Judge from previous work is that, while we also enable dynamic evaluation through simulated interaction, unlike prior work that evaluates the utterances of simulated interactions (Zhou et al., 2023d; Wu et al., 2025; Shaikh et al., 2024), we focus on simulating human-like authentic feedback for the agent systems based on the user’s emotional changes, inner thoughts, and utterances, achieved through the principled reasoning process of the Sentient Agent. Additionally, our framework is highly adaptable to various evaluation sub-scenarios for assessing LLM agents’ higher-order social cognition.
6 Conclusion
This work presents Sentient Agent as a Judge, a novel framework for evaluating the higher-order social-cognitive abilities of LLMs in emotionally complex dialogues. By grounding assessments in simulated users endowed with personas, goals, and adaptive emotional feedback, our approach offers a scalable and interpretable benchmark that more accurately reflects real-world expectations of social interaction. Through extensive experiments with 18 foundation models, we demonstrate that Sentient emotion scores capture meaningful distinctions in empathy quality and conversational attunement, aligning strongly with both internal user thoughts and discourse evaluations. The Sentient Leaderboard and our Social Cognition Coordinate chart reveal that mastery of social reasoning lags behind linguistic competence.
Looking ahead, we plan to expand the scenario library to cover negotiation, deception detection and multicultural contexts, and investigate training curricula that directly optimize for Sentient feedback. We hope Sage will serve both researchers and practitioners as a rigorous yard‑stick and a catalyst for building language agents that are not only coherent and knowledgeable, but also genuinely human‑sensitive.
Appendix A Definition of Strategies
When faced with the task of supporting Sentient Agents, each LLM applies its unique response style, which typically involves a mix of question asking, comforting, and providing suggestions. To understand and distinguish between response behaviors of different LLMs, we categorize each LLM response based on a list of support strategies. Our construction of support strategies is inspired by Liu et al. (2021), although we modify and split their main groups of strategies into fine-grained strategies. A list of available strategies is in Table 3.
To analyze the support strategies used by an LLM, we prompted DeepSeek-V3 to act as a judge, evaluating each round of the model output to identify all support strategies involved. We then aggregated the strategy statistics across all rounds of conversations, outputing the proportion of rounds each strategy is used. The resulting statistics are then used as one contributing factor in our analysis of the Social Cognition Coordinate system.
| Group | Strategy |
| (A) Question | (A-1) Information-seeking questions |
| (A-2) Asking about the client’s mental state | |
| (A-3) Asking the client whether a solution has been attempted | |
| (A-4) Reflective questions about the client’s views | |
| (A-5) Rhetorical questions | |
| (B) Emotional Empathy | (B-1) Surface-level empathy |
| (B-2) Providing empathy via restating the client’s problem | |
| (B-3) Deeper empathy to understand the client’s hidden intention | |
| (C) Self-Disclosure | (C-1) Self-disclosure that provides agreement with the client’s view |
| (C-2) Self-disclosure that introduces the supporter’s own story | |
| (D) Emotional Comfort | (D-1) Providing comforting words to the client |
| (D-2) Expressing willingness to hear the client’s thoughts | |
| (D-3) Helping the client to vent negative feelings | |
| (E) Affirmation and Reassurance | (E-1) Praising the client’s qualities |
| (E-2) Praising the client’s positive thoughts | |
| (E-3) Praising the client’s actions | |
| (E-4) Providing accompaniment and support | |
| (F) Providing Suggestions | (F-1) Analysis of the client’s issue |
| (F-2) Suggestions for emotional relief | |
| (F-3) Suggestions for seeking psychological counseling | |
| (F-4) General advice for solving client’s issue | |
| (F-5) Advice specific to the client’s situation | |
| (G) Information | (G-1) Information related to emotional support |
| (G-2) Information related to problem-solving suggestions |
Appendix B Factors of Success and Failure
To determine the social cognition profile of each LLM, we analyze the factors contributing to their success (emotion score 100) or failure (emotion score 10) in the interaction during benchmarking. Specifically, for each LLM in our benchmark, we randomly select 5 successful and 5 failed cases from Sentient Agents exhibiting both emotional and rational intentions, resulting in 20 cases per model. We then prompt Gemini2.5-Pro to analyze the underlying reasons for each model’s success or failure in social cognition during interaction. Subsequently, we present these analytical results to Gemini2.5-Pro again, asking it to summarize the distinguishing characteristics of each LLM’s approach in the interaction with Sentient Agents, with particular attention to aspects such as social distance, professionalism, and personality. In Figure 10, we present example outputs of the analysis for four representative models examined in the case study (§4.4), which offer a different perspective on interpreting the cases from Gemini2.5-Pro’s viewpoint.